NotesCambridgeO LevelMathematics
45.9 KB · 0 downloads
Shared by Nobold ZIMFundi team · 4 days ago
P(event) = number of favourable outcomes ÷ total equally likely outcomes; 0 ≤ P ≤ 1; P(not A) = 1 − P(A). Relative frequency estimates probability from experiments; expected frequency = P × number of trials.
Comments (0)
No comments yet.
Sign in or create a free account to join in.